Joe drove to work on Monday at 45 mph and arrived 1 minute early. On Tuesday he drove at 40 mph and arrived 1 minute late.
a) How far does Joe live from work?
b) What speed does he normally travel?
c) How fast should he go to arrive 5 minutes early?

Answer by checkley75(3666)

d=rt
45(t-1/60)=40(t+1/60)
45t-45/60=40t+45/60
45t-40t=90/60
5t=1.5
t=1.5/5
t=3 hours is his normal driving time.
a)45(3-1/60)=135-45/60=135-.75=134.25 miles.
b)(45-40)/2=85/2=42.5 mph.
c) 134.25=r(3-5/60)
134.25=3r-15/60
134.25+.25=3r
134.50=3r
r=134.5/3
r=44.8333 mph is his speed in order to arive 5 minutes early.